OpenStudy (naveenbhatia1312):

Find all solutions in the interval [0, 2π). sin^2x - cos^2x = 0

OpenStudy (anonymous):

hhaha funny here are the (6 points)

OpenStudy (danjs):

i gave you the biggest clue on your last post \[\sin^2(a) - (1 - \sin^2(a)) = 0\] sin^2(a) = 1/2 so \[\sin(a) = \pm \frac{ 1 }{ \sqrt{2} }\]

OpenStudy (danjs):

that is the same as \[\sin(a) = \pm \frac{ \sqrt{2} }{ 2 }\] each point on the unit circle is \[(x,y) = (\cos(a), \sin(a))\] so the y coordinate is the sin(a) value look where that is root2/2 or -root(2)/2

OpenStudy (danjs):

4 points, 4 angles

OpenStudy (naveenbhatia1312):

this is confusing

OpenStudy (danjs):

need to practice the easier stuff first

OpenStudy (naveenbhatia1312):

whats the answer?

OpenStudy (danjs):

I used the identity sin^2 + cos^2 = 1 the other is just algebra to solve for sin(a) sin(a) is that value for certain angles on the unit circle...

Parth (parthkohli):

Just another fun solution. It's akin to saying that \(-\cos(2x) = 0\) so \(\cos(2x) = 0\).

OpenStudy (danjs):

angles on here where y coordinate is that value, there are 4

OpenStudy (danjs):

goodluck

OpenStudy (naveenbhatia1312):

@ParthKohli what is the answer?

Parth (parthkohli):

How would you solve \(\cos(2x) = 0\)? In general when is cosine zero?

OpenStudy (naveenbhatia1312):

pi/2 or 3pi/2

OpenStudy (naveenbhatia1312):

@ParthKohli